# Making Teams and Influencing Agents: Efficiently Coordinating Decision Trees for Interpretable Multi-Agent Reinforcement Learning

### v1

@@ 1 Introduction
Over the past decade, multi-agent reinforcement learning (MARL) algorithms have achieved state-of-the-art performance in various challenging board and video games (Silver et al. 2016; Vinyals et al. 2019). They have also found success in other sequential decision-making scenarios based on critical real-world domains, including robotics (Orr and Dutta 2023), cybersecurity (Panfili et al. 2018), and traffic signal control (TSC) (Chen et al. 2020). However, the real-world applicability of these algorithms is hampered by two key challenges. First, the deep neural network (NN) architectures that are needed to achieve good performance have thousands to millions of parameters. Second, the behavior of RL agents is difficult to predict and verify due to its dependency on complex state spaces and long time horizons. Thus, human stakeholders understand and trust RL agents less than their simpler counterparts, even if RL yields superior performance (Siu et al. 2021). Conversely, more interpretable representations of policies can help stakeholders to build appropriate levels of trust in RL agents (Druce, Harradon, and Tittle 2021; Zhang, Liao, and Bellamy 2020).
@@
In applications where the safety and verifiability of RL policies is critical (Gilbert et al. 2023; Jayawardana, Landler, and Wu 2021), such as TSC, users may deploy interpretable surrogate policies in place of expert NN policies. Such surrogate policies should be performant — capable of achieving high returns. In MARL, coordinating the training of surrogates is critical for performance: if multiple surrogates are deployed simultaneously, they cannot assume that they are interacting with performant experts, as their performance may be influenced by suboptimal behaviour from others.
@@
At the same time, surrogate policies should be computationally efficient — it should be possible to generate them with minimal environment interactions and runtime. Frequent real-world environment interactions are impractical in domains such as robotics (Finn et al. 2017) and TSC (Zang et al. 2020), and simulators remain computationally intensive if they need to be of sufficient fidelity to guarantee real-world transferrability (Chen et al. 2023). In human-in-the-loop frameworks where users provide oversight to correct undesirable policy behaviour (Mandel et al. 2017), the ability to iterate on surrogate policies in a rapid fashion is also critical (Wu et al. 2023). The desiderata of performance and computational efficiency exist in tension: more complex models capable of stronger performance and coordination capabilities are less efficient (Milani et al. 2024).
@@
Decision trees (DTs) have been considered an attractive model structure for interpretable single-agent RL due to their comprehensibility (Silva et al. 2020). They lie at the core of the imitation learning framework VIPER (Bastani, Pu, and Solar-Lezama 2018), which has been applied to distill NN-based RL policies into DTs in domains such as TSC (Jayawardana, Landler, and Wu 2021; Zhu, Yin, and Chen 2022), autonomous vehicles (Schmidt et al. 2021), and robotics (Roth et al. 2023). However, generalizing VIPER to the MARL setting is challenging. Prior work (Milani et al. 2022) has introduced two multi-agent VIPER algorithms, IVIPER and MAVIPER. IVIPER fails to coordinate the training of DTs, thus sacrificing performance; meanwhile, MAVIPER trains DTs in a coordinated but computationally inefficient manner. This makes DT surrogates generated by these algorithms impractical for deployment.
@@
To this end, we introduce HYDRAVIPER, an efficient method to extract coordinated DT policies for cooperative MARL. Our method makes three key algorithmic contributions: (1) HYDRAVIPER coordinates agent training by jointly resampling the training dataset for each team of cooperative agents. (2) When interacting with the environment to collect a training dataset, HYDRAVIPER adaptively collects critical trajectories closer to convergence. (3) When interacting with the environment for evaluation, HYDRAVIPER uses a multi-armed bandit-based evaluation strategy to identify promising sets of trained surrogates. Experiments on various environments demonstrate that HYDRAVIPER achieves our goal of balancing performance and computational efficiency. HYDRAVIPER also improves the practical applicability of DT-based interpretable MARL policies: users can exchange training time for performance by altering the algorithm’s environment interaction budgets, but its performance remains optimal at different budget levels. Lastly, HYDRAVIPER’s efficiency on large environments can be improved while maintaining coordination by dividing the agent set into mutually influential teams.
@@ 2 Related Work
Past methods for interpretable MARL have focused on using feature importance measures to construct saliency maps (Iqbal and Sha 2019; Heuillet, Couthouis, and Díaz-Rodríguez 2022; Liu, Zhu, and Chen 2023; Motokawa and Sugawara 2023), logical structures (Kazhdan, Shams, and Lio 2020; Wang et al. 2021; Ji, Li, and Xiao 2023), and domain concepts (Zabounidis et al. 2023). Each of these categories of methods has limitations. Feature importances and saliency maps are visua lly clear, but only highlight aspects of the state space without showing how policies use them. Policies based on logical rules and concepts allow users to align the execution of these policies with domain knowledge, but they require extensive feature engineering. By contrast, we learn simple policy representations grounded directly in the environment feature space.
@@
Relative to deep NNs, shallow DT policy representations are intrinsically (Molnar 2019) and empirically (Silva et al. 2020) more comprehensible. One line of work in DT-based RL directly trains DT policies (Silva et al. 2020; Topin et al. 2020; Crespi et al. 2023; Liu et al. 2024) using relaxations amenable to direct optimization. However, these methods suffer from training instability and performance degradation. Another line of work follows the VIPER framework (Bastani, Pu, and Solar-Lezama 2018), in which a surrogate DT is trained through imitation learning of a performant expert. Although VIPER has achieved success in single-agent settings (Schmidt et al. 2021; Roth et al. 2023; Jayawardana, Landler, and Wu 2021; Zhu, Yin, and Chen 2022), only two VIPER-based algorithms exist for the multi-agent setting: IVIPER and MAVIPER (Milani et al. 2022). IVIPER independently trains DTs for each agent in a decentralised manner, enjoying computational efficiency at the cost of performance due to lack of coordination. MAVIPER jointly trains DTs in a centralised manner to achieve coordination, but it suffers from computational inefficiency. Neither algorithm balances performance and computational efficiency.
@@ 3 Background
Figure 1: (L) Imitation learning in traffic signal control, where a decision tree must be learned to imitate the RL-based signal control policy of each intersection’s controller agent (the seven-intersection Ingolstadt corridor TSC environment from Section 5.1 is shown, with intersections numbered). (R) The HYDRAVIPER framework, in which DT students are trained independently using a jointly resampled dataset of trajectories collected from an environment and relabeled by an NN expert.
@@
We model agent cooperation as a team-based Markov game. A Markov game for $N$ agents consists of a set of states $\mathcal{S}$ with initial state distribution $\rho:\mathcal{S}\rightarrow[0,1]$, and a set of actions $\mathcal{A}_{1},...,\mathcal{A}_{N}$ and observations $\mathcal{O}_{1},...,\mathcal{O}_{N}$ (consisting of features correlated with the state) for each agent $i$. The agent set is partitioned into disjoint teams $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\subseteq[N]$. Each agent chooses actions according to a policy $\pi_{i}:\mathcal{O}_{i}\rightarrow\mathcal{A}_{i}$. After agents simultaneously execute actions $\mathbf{a}$, the environment produces the next state based on the state transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\ldots\times\mathcal{A}_{N}\rightarrow\mathcal{S}$, a joint reward for each team $\mathcal{T}_{\ell}$ based on $R_{\ell}:\mathcal{S}\times\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\rightarrow\mathbb{R}$, and private observations $O_{i}:\mathcal{S}\rightarrow\mathcal{O}_{i}$. Each agent maximises its team’s return $\mathcal{R}_{\ell}=\sum_{t=0}^{T}\gamma^{t}R_{\ell}(s_{t},\pi(\mathbf{o}_{t}))$ over $T$ timesteps, where $\gamma$ is a discount factor weighting the importance of future rewards.
@@
    Collects $K$ new rollouts $\{\mathbf{o},\hat{\pi}^{m-1}(\mathbf{o})\}$ using the previous students from iteration $m-1$ (where $\hat{\pi}^{0}:=\pi^{*}$)
@@
After $M$ iterations, VIPER (5) selects a student through validation on an additional set of rollouts. Ross, Gordon, and Bagnell (2011) showed that such a procedure is guaranteed to find a student that is performant on the distribution of states that it induces.
@@ 4 HYDRAVIPER
In this section, we present HYDRAVIPER (Algorithm 1), our algorithm for performant and efficient interpretable MARL. As shown in Figure 1, HYDRAVIPER builds on the DAGGER and VIPER frameworks by iteratively collecting data from environment rollouts to train DT policies. HYDRAVIPER first (1) partitions agents into clusters for scalability (line 4). Next, in each of $M$ iterations, HYDRAVIPER: (2) collects a dataset of rollouts from the environment, using an adaptive procedure (lines 6–7); (3) resamples the dataset to prioritise learning the correct actions in critical states, using team-based $Q$-values (lines 9–10); (4) and trains DTs based on these datasets (lines 11–12). After completing all $M$ training iterations, HYDRAVIPER (5) identifies the best-performing student for each agent, using a multi-armed bandit algorithm, and returns them as a policy profile (lines 13–14). We now describe each of these algorithm components in detail.
@@
14:         $\mathcal{D}_{\ell}^{\prime},K_{\mathrm{drop}}^{\prime}\leftarrow$ C-Q$(\mathcal{D}_{\ell},\mathcal{T}_{\ell},\pi^{*},Q^{\pi^{*}},\epsilon)$
@@
20:     $\hat{\pi}_{i},\forall i\in\mathcal{T}_{\ell}\leftarrow$ VR-UCB($(\{\hat{\pi}^{m}_{\ell}\}_{m=1}^{M},\mathcal{T}_{\ell},B_{\mathrm{valid}},c)$
@@ 4.1 Dataset Resampling: Centralised-Q Weighting
1:Input: Dataset $\mathcal{D}_{\ell}$, teams $\mathcal{T}_{\ell}$, experts $\pi^{*}$, expert $Q$-functions $Q^{\pi^{*}}$, threshold $\epsilon$
@@
VIPER-based algorithms include a dataset resampling step (Algorithm 1, lines 9–10) so that students can focus their learning on more critical states. At a high level, they construct a training dataset by computing sampling weights over the aggregated dataset of environment rollouts, typically using some notion of value based on the expert $Q$-functions. Measuring value is straightforward in the single-agent setting, but — as we have mentioned — a key obstacle in multi-agent learning is efficient coordination among agents. To address this challenge, HYDRAVIPER induces coordination in the resampling step using a team-based notion of value (Algorithm 2), but trains DTs independently for each agent.
Specifically, HYDRAVIPER resamples the dataset for DT construction based on weights $p_{\ell k}$, which compute the relative importance of each sample for each team of agents $\mathcal{T}_{\ell}$ (Algorithm 2, line 1). Prior work computes this importance based on individual $Q$-functions, meaning that each agent must maintain its own dataset and induce coordination through (typically computationally expensive) joint training procedures. By contrast, we propose an intuitive change: HYDRAVIPER uses the mean of the expert $Q$-functions within each team of coordinated agents, $\bar{Q}^{\pi_{\ell}^{*}}:=\frac{1}{|\mathcal{T}_{\ell}|}\sum_{j\in\mathcal{T}_{\ell}}Q^{\pi_{j}^{*}}$, so as to prioritise samples according to their value to the team. Then, we compute the weights as the difference in value between the optimal joint team action and the worst-case joint team action. Intuitively, highly-weighted samples are those where coordinating on joint actions matters for performance. The weights are defined as:
@@
For further gains in sample efficiency, HYDRAVIPER does not compute $p_{\ell k}$ by enumerating joint actions over all agents present in the environment. Instead, it only enumerates possible joint actions $\mathbf{a}_{\ell}$ over the team and uses expert actions $\pi_{-\ell}^{*}(\mathbf{o}_{-\ell})$ for opponent agents. This novel resampling procedure eliminates the need for per-agent datasets in IVIPER and MAVIPER, allowing agents to prioritise the same critical states without computationally expensive joint training.
HYDRAVIPER uses the jointly sampled dataset to independently train DTs for each agent $i$ (Algorithm 1, lines 8–9). Each DT $\hat{\pi}_{i}$ uses individual observations $o_{i}$ to fit $\pi_{i}^{*}$’s actions in the dataset. Modifying the input dataset rather than the training procedure provides HYDRAVIPER with flexibility in the choice of DT learning algorithm. More advanced models such as random forests or mixtures of DTs (Vasić et al. 2022) can be incorporated to improve performance.
@@ 4.2 Training Rollouts: Adaptive Budget Allocation
Thus far, we have assumed that HYDRAVIPER has access to a dataset of observation-action pairs for training. To collect this dataset, HYDRAVIPER follows the DAGGER-style iterative procedure of collecting a dataset at each iteration $m$ by rolling out the current student policies $\hat{\pi}^{m-1}$ (Algorithm 1, line 6–7). The next set of students are trained on the aggregate of all collected datasets, therefore building up the set of inputs likely to be encountered by the student policies during execution. However, collecting training rollouts is computationally expensive. Prior work employs an inefficient static allocation strategy which uniformly performs $K_{\textrm{train}}$ rollouts in each iteration. This strategy is problematic because the students are far from convergence early in training, so the distribution of trajectories collected earlier in training potentially diverges from those that converged students would encounter. HYDRAVIPER addresses this challenge through an adaptive rollout strategy that dynamically allocates the training budget at each iteration and prioritises critical states encountered later in training.
@@
Specifically, we treat the first iteration as a warm-up period, in which the experts collect a predefined number of $K_{\mathrm{train}}$ rollouts (Algorithm 3, lines 3–5). This leads to an initial dataset of $T\cdot K_{\mathrm{train}}$ observation-action pairs. Each team $\mathcal{T}_{\ell}$ discards non-critical samples from its dataset (Algorithm 2, line 2), i.e. those where the range in the $Q$-value is at most a predefined threshold $\epsilon$. With the goal of efficiency in mind, HYDRAVIPER computes the minimum number of such discarded samples across all teams of cooperative agents. This then determines the minimum number of rollouts required to collect at least this many samples in the next iteration. As a result, the expected number of dropped rollouts, and therefore the budget for the next iteration, is:
$K_{\mathrm{drop}}=\min_{\ell}\left\lceil{\frac{1}{T}|\{(\mathbf{o}_{k},\mathbf{a}_{k})\in\mathcal{D}_{\ell}\mid p_{\ell k}\leq\epsilon\}|}\right\rceil.$
Over the remaining $M-1$ iterations, HYDRAVIPER continues collecting rollouts using students until it exhausts its total budget of $B_{\mathrm{train}}$ training rollouts (Algorithm 3, line 1). Choosing different rollout budgets allows performance and efficiency to be traded off. A higher budget is likely to lead to superior performance, as more rollouts will be collected from students closer to convergence before the budget is exhausted, but it also requires more computation time.

### v2

@@ 1 Introduction
Over the past decade, multi-agent reinforcement learning (MARL) algorithms have achieved state-of-the-art performance in various challenging board and video games (Silver et al. 2016; Vinyals et al. 2019). They have also found success in other sequential decision-making scenarios based on critical real-world domains, including robotics (Orr and Dutta 2023), cybersecurity (Panfili et al. 2018), and traffic signal control (TSC) (Chen et al. 2020). However, the real-world applicability of these algorithms is hampered by two key challenges. First, the deep neural network (NN) architectures that are needed to achieve good performance have thousands to millions of parameters. Second, the behaviour of RL agents is difficult to predict and verify due to its dependence on complex state spaces and long time horizons. Thus, human stakeholders understand and trust RL agents less than their simpler counterparts, even if RL yields superior performance (Siu et al. 2021). Conversely, more interpretable representations of policies can help stakeholders build appropriate levels of trust in RL agents (Druce, Harradon, and Tittle 2021; Zhang, Liao, and Bellamy 2020).
@@
In applications where the safety and verifiability of RL policies is critical (Gilbert et al. 2023; Jayawardana, Landler, and Wu 2021), such as TSC, users may deploy interpretable surrogate policies instead of expert NN policies. Such surrogate policies should be performant — capable of achieving high returns. In MARL, coordinating the training of surrogates is critical for performance: if multiple surrogates are deployed simultaneously, they cannot assume that they are interacting with performant experts, as their performance may be influenced by other agents’ suboptimal behaviour.
@@
At the same time, surrogate policies should be computationally efficient; it should be possible to generate them with minimal environment interactions and runtime. Frequent interactions with the environment are impractical in domains such as robotics (Finn et al. 2017) and TSC (Zang et al. 2020), and simulators that have sufficient fidelity for real-world transferrability (Chen et al. 2023) are computationally intensive. In human-in-the-loop frameworks where users provide oversight to correct undesirable policy behaviour (Mandel et al. 2017), the ability to quickly iterate on surrogate policies is also critical (Wu et al. 2023). More complex models capable of stronger performance and coordination capabilities are less efficient (Milani et al. 2024).
@@
Decision trees (DTs) are an attractive model class for interpretable RL due to their comprehensibility (Silva et al. 2020). They also enable the design of responsible AI systems, as their branching rules can be easily verified and constrained by human experts or automated processes (Blockeel et al. 2023). DTs lie at the core of the imitation learning framework VIPER (Bastani, Pu, and Solar-Lezama 2018), which has been applied to distil NN-based RL policies into DTs in domains such as TSC (Jayawardana, Landler, and Wu 2021), autonomous vehicles (Schmidt et al. 2021), and robotics (Roth et al. 2023). However, generalising VIPER to the MARL setting is challenging. Past work (Milani et al. 2022) introduced two multi-agent VIPER algorithms, IVIPER and MAVIPER, which are both impractical for deployment. IVIPER fails to coordinate the training of DTs, thus sacrificing performance; MAVIPER trains DTs in a coordinated but computationally inefficient manner.
@@
To this end, we introduce HYDRAVIPER [^1] Code: https://github.com/lythronaxargestes/hydraviper-public, an efficient method to extract coordinated DT policies for cooperative MARL. Our method makes three key algorithmic contributions: (1) HYDRAVIPER coordinates agent training by jointly resampling the training dataset for each team of cooperative agents. (2) When interacting with the environment to collect a training dataset, HYDRAVIPER adaptively collects critical trajectories closer to convergence. (3) When interacting with the environment for evaluation, HYDRAVIPER uses a multi-armed bandit-based evaluation strategy to identify promising sets of trained surrogates. Experiments demonstrate that HYDRAVIPER achieves our goal of balancing performance and computational efficiency. HYDRAVIPER also improves the applicability of DT-based interpretable MARL policies: users can exchange training time for performance by altering its environment interaction budgets, but its performance remains optimal at different budget levels. Lastly, HYDRAVIPER’s efficiency on large environments can be improved while maintaining coordination by dividing the agent set into mutually influential teams.
@@ 2 Related Work
Past methods for interpretable MARL have focused on using feature importance measures to construct saliency maps (Iqbal and Sha 2019; Heuillet, Couthouis, and Díaz-Rodríguez 2022; Liu, Zhu, and Chen 2023; Motokawa and Sugawara 2023), building logical structures (Kazhdan, Shams, and Lio 2020; Wang et al. 2021; Ji, Li, and Xiao 2023), and defining domain concepts (Zabounidis et al. 2023). Each of these categories of methods has limitations. Feature importances and saliency maps are visually clear, but only highlight aspects of the state space without showing how policies use them. Policies based on logical rules and concepts allow users to align the execution of these policies with domain knowledge, but require extensive feature engineering. By contrast, we learn simple policy representations grounded directly in the environment feature space.
@@
Relative to deep NNs, shallow DT policy representations are intrinsically (Molnar 2019) and empirically (Silva et al. 2020) more comprehensible. One line of work in DT-based RL directly trains DT policies (Silva et al. 2020; Topin et al. 2020; Crespi et al. 2023; Liu et al. 2025) using relaxations amenable to direct optimisation. However, these methods suffer from training instability and performance degradation. Another line of work follows the VIPER framework (Bastani, Pu, and Solar-Lezama 2018), in which a surrogate DT is trained by imitation learning of a performant expert. Although VIPER has achieved success in single-agent settings (Schmidt et al. 2021; Roth et al. 2023; Jayawardana, Landler, and Wu 2021; Zhu, Yin, and Chen 2022), only two VIPER-based algorithms exist for multi-agent settings: IVIPER and MAVIPER (Milani et al. 2022). IVIPER independently trains DTs for each agent in a decentralised manner, enjoying computational efficiency at the cost of performance due to a lack of coordination. MAVIPER jointly trains DTs in a centralised manner to achieve coordination, but suffers from computational inefficiency. Neither algorithm balances performance and computational efficiency.
@@ 3 Background
Figure 1: (L) Imitation learning in traffic signal control, where a decision tree must be learnt to imitate the RL-based policy of each intersection’s signal controller agent (the seven-intersection Ingolstadt corridor TSC environment from Section 5.1 is shown, with intersections numbered). (R) The HYDRAVIPER framework, in which DT students are trained independently using a jointly resampled dataset of environment trajectories and relabelled by an NN expert.
@@
We model agent cooperation as a team-based Markov game. A Markov game for $N$ agents consists of a set of states $\mathcal{S}$ with initial state distribution $\rho:\mathcal{S}\rightarrow[0,1]$, and sets of actions $\mathcal{A}_{1},...,\mathcal{A}_{N}$ and observations $\mathcal{O}_{1},...,\mathcal{O}_{N}$ (consisting of features correlated with the state) for each agent $i$. The agent set is partitioned into disjoint teams $\mathcal{T}_{1},\ldots,\mathcal{T}_{L}\subseteq[N]$. Each agent chooses actions according to a policy $\pi_{i}:\mathcal{O}_{i}\rightarrow\mathcal{A}_{i}$. After agents simultaneously execute actions $\mathbf{a}$, the environment produces the next state based on the state transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\ldots\times\mathcal{A}_{N}\rightarrow\mathcal{S}$, a joint reward for each team $\mathcal{T}_{\ell}$ based on $R_{\ell}:\mathcal{S}\times\prod_{i\in\mathcal{T}_{\ell}}\mathcal{A}_{i}\rightarrow\mathbb{R}$, and private observations $O_{i}:\mathcal{S}\rightarrow\mathcal{O}_{i}$. Each agent maximises its team’s return $\mathcal{R}_{\ell}=\sum_{t=0}^{T}\gamma^{t}R_{\ell}(s_{t},\pi(\mathbf{o}_{t}))$ over $T$ timesteps, where $\gamma$ is a discount factor weighting the importance of future rewards.
@@
    Collects $K$ new rollouts $\{\mathbf{o},\hat{\pi}^{m-1}(\mathbf{o})\}$ using the previous students from iteration m1m-1 (where $\hat{\pi}^{0}:=\pi^{*}$)
@@
After $M$ iterations, VIPER (5) selects a student through validation on an additional set of rollouts. Ross, Gordon, and Bagnell (2011) showed that such a procedure is guaranteed to find a student which is performant on the distribution of states that it induces.
@@ 4 HYDRAVIPER
In this section, we present HYDRAVIPER (Algorithm 1), our algorithm for performant and efficient interpretable MARL. As shown in Figure 1, HYDRAVIPER builds on the DAGGER and VIPER frameworks by iteratively collecting data from environment rollouts to train DT policies. HYDRAVIPER first (1) partitions agents into clusters for scalability (line 4). Next, in each of $M$ iterations, HYDRAVIPER: (2) collects a dataset of rollouts from the environment, using an adaptive procedure (lines 6–7); (3) resamples the dataset to prioritise learning the correct actions in critical states, using team-based $Q$-values (lines 9–10); (4) and trains DTs based on these datasets (lines 11–12). After it completes all $M$ training iterations, HYDRAVIPER (5) identifies the best-performing student for each agent, using a multi-armed bandit algorithm, and returns them as a policy profile (lines 13–14). Now, we describe each of these algorithm components in detail.
@@
[…]
@@
[…]
@@ 4.1 Dataset Resampling: Centralised-Q Weighting
1:Input: Team dataset $\mathcal{D}_{\ell}$, team $\mathcal{T}_{\ell}$, experts $\pi^{*}$, expert $Q$-functions $Q^{\pi^{*}}$, threshold $\epsilon$
@@
VIPER-based algorithms include a dataset resampling step (Algorithm 1, lines 9–10) so that students can focus their learning on more critical states. At a high level, they construct a training dataset by computing sample weights on the aggregated dataset of environment rollouts, typically using some notion of value based on the expert $Q$-functions. Measuring value is straightforward in the single-agent setting, but — as we have mentioned — a key obstacle in multi-agent learning is efficient coordination among agents. To address this challenge, HYDRAVIPER induces coordination in the resampling step using a team-based notion of value (Algorithm 2), but trains DTs independently for each agent.
Specifically, HYDRAVIPER resamples the dataset $\mathcal{D}$ for DT construction based on weights $p_{\ell k}$, which represent the relative importance of each sample for each team of agents $\mathcal{T}_{\ell}$ (Algorithm 2, line 1). Past work computed this importance based on individual $Q$-functions, meaning that each agent must maintain its own dataset $\mathcal{D}_{i}$ and induce coordination through (typically computationally expensive) joint training procedures. By contrast, we propose an intuitive change: HYDRAVIPER uses the mean of the expert $Q$-functions within each team of coordinated agents, $\bar{Q}^{\pi_{\ell}^{*}}:=\frac{1}{|\mathcal{T}_{\ell}|}\sum_{j\in\mathcal{T}_{\ell}}Q^{\pi_{j}^{*}}$, to prioritise samples according to their value to the team. Then, we compute the weights as the difference in value between the optimal joint team action and the worst-case joint team action. Intuitively, highly-weighted samples are those where coordinating on joint actions matters for performance. The weights are defined as:
@@
For further gains in sample efficiency, HYDRAVIPER does not compute $p_{\ell k}$ by enumerating joint actions over all agents in the environment. Instead, it only enumerates possible joint actions $\mathbf{a}_{\ell}$ over the team and uses expert actions $\pi_{-\ell}^{*}(\mathbf{o}_{-\ell})$ for the opponent agents. This novel resampling procedure eliminates the need for per-agent datasets in IVIPER and MAVIPER, allowing agents to prioritise the same critical states without computationally expensive joint training.
HYDRAVIPER uses each team’s jointly sampled dataset $\mathcal{D}_{\ell}$ to independently train DTs for each agent $i$ (Algorithm 1, lines 8–9). The DT $\hat{\pi}_{i}$ uses individual observations $o_{i}$ to fit $\pi_{i}^{*}$’s actions in the dataset. Modifying the input dataset rather than the training procedure provides HYDRAVIPER with flexibility in the choice of DT learning algorithm. We use CART (Breiman et al. 1984), but more advanced models such as random forests or mixtures of DTs (Vasić et al. 2022) can also be used to improve performance.
@@ 4.2 Training Rollouts: Adaptive Budget Allocation
Thus far, we have assumed that HYDRAVIPER has access to a dataset of observation-action pairs for training. To collect this dataset, HYDRAVIPER follows the DAGGER-style iterative procedure of collecting a dataset at each iteration $m$ by rolling out the current student policies $\hat{\pi}^{m-1}$ (Algorithm 1, line 6–7). The next set of students are trained on the aggregate of all collected datasets, therefore building up the set of inputs likely to be encountered by the student policies during execution. However, collecting training rollouts is computationally expensive. Past work has employed an inefficient static allocation strategy that uniformly performs $K_{\textrm{train}}$ rollouts in each iteration. This strategy is problematic because the students are far from convergence early in training, so the distribution of trajectories collected earlier in training potentially diverges from those that converged students would encounter. HYDRAVIPER addresses this challenge through an adaptive rollout strategy that dynamically allocates the training budget at each iteration and prioritises critical states encountered later in training.
@@
Specifically, we treat the first iteration as a warm-up period, in which the experts collect a predefined number of $K_{\mathrm{train}}$ rollouts (Algorithm 3, lines 3–5). This leads to an initial dataset of $T\cdot K_{\mathrm{train}}$ observation-action pairs. Each team $\mathcal{T}_{\ell}$ discards non-critical samples from its dataset (Algorithm 2, line 2), i.e. those where the range in the $Q$-value is at most a predefined threshold $\epsilon$. With the goal of efficiency in mind, HYDRAVIPER computes the minimum number of such discarded samples across all teams of cooperative agents. This then determines the minimum number of rollouts required to collect at least this many samples in the next iteration. The expected number of dropped rollouts, and therefore the budget for the next iteration, is:
$K_{\mathrm{drop}}=\min_{\ell}\left\lceil{\frac{1}{T}\left|\{(\mathbf{o}_{k},\mathbf{a}_{k})\in\mathcal{D}_{\ell}\mid p_{\ell k}\leq\epsilon\}\right|}\right\rceil.$
During the remaining M1M-1 iterations, HYDRAVIPER continues to collect rollouts using students until it exhausts its total budget of $B_{\mathrm{train}}$ training rollouts (Algorithm 3, line 1). Choosing different rollout budgets allows performance and efficiency to be traded off. A higher budget is likely to lead to superior performance, as more rollouts will be collected from students closer to convergence before the budget is exhausted, but it also requires more computation time.